2) A population of salamanders from Georgia what genotyped at a single autosomal allozyme locus with two alleles (T and L) as follows: Genotype TT TL LL Total Number 375 140 18 533

a) What are the genotype frequencies of this population?

b) What are the allele frequencies of this population?

c) Assuming this population is in Hardy-Weinberg equilibrium, predict the genotype frequencies based on the allele frequencies.
Answer #1

According to Hardy-Weinberg equilibrium law

B. P2 + 2Pq + q2 = 1 and P + q = 1

The recessive allele frequency q2 = 18/533 = 0.033

                 Take square root q = 0.18

P + q = 1 ; P = 1 - q = 1 - 0.18 = 0.82

A. Genotype frequencies P2 = 0.82 X 0.82 = 0.67

Genotype frequencies 2Pq = 2 X 0.82 X 0.18 = 0.29

Genotype frequencies q2 = 0.18 X 0.18 = 0.033